# Beamfocusing Optimization for Near-Field Wideband Multi-User Communications

### v4

@@ Abstract
A near-field wideband communication system is investigated in which a base station (BS) employs an extra-large scale antenna array (ELAA) to serve multiple users in its near-field region. To facilitate near-field multi-user beamforming and mitigate the spatial wideband effect, the BS employs a hybrid beamforming architecture based on true-time delayers (TTDs). In addition to the conventional fully-connected TTD-based hybrid beamforming architecture, a new sub-connected architecture is proposed to improve energy efficiency and reduce hardware requirements. Two wideband beamforming optimization approaches are proposed to maximize spectral efficiency for both architectures. 1) Fully-digital approximation (FDA) approach based on full channel state information (CSI): In this method, the TTD-based hybrid beamformer is optimized by the block-coordinate descent and penalty method to approximate the optimal digital beamformer. This approach ensures convergence to the stationary point of the spectral efficiency maximization problem. 2) Heuristic two-stage (HTS) approach based on partial CSI: In this approach, a piecewise-near-field approximation of near-field channels is first proposed to facilitate the design of TTD-based analog beamformers based on the outcomes of near-field beam training. Subsequently, the low-dimensional digital beamformer is optimized using knowledge of the low-dimensional equivalent channels, resulting in reduced computational complexity and channel estimation complexity. Our numerical results show that 1) the proposed approach effectively eliminates the spatial wideband effect, and 2) the proposed sub-connected architecture is more energy efficient and has fewer hardware constraints on the TTD and system bandwidth compared to the fully-connected architecture.
@@
Beamfocusing, near-field communications, spatial wideband effect, true-time delayers, wideband communications
@@ I Introduction
Near-field effect
Spatial wideband effect
Multiple communication users
The employment of TTDs
@@ I-B Motivations and Contributions
Against the above background, the TTD-based hybrid beamforming architecture is a promising solution for the wideband beamforming design in either far-field or near-field systems. However, it is notable that most existing research in this area has predominantly focused on single-user systems [13, 15, 16, 22, 17, 18, 20, 21]. In such contexts, beamforming design is typically oriented towards maximizing array gain along the direction or location of the user. This approach, while effective for single-user scenarios, is flawed in multi-user systems where both array gain and inter-user interference must be carefully balanced. This requirement introduces a complex design challenge, particularly in the context of TTD-based hybrid beamforming. To the best of the authors’ knowledge, this is the first attempt to directly address the design problem of TTD-based hybrid beamforming in multiuser systems by considering both unit modulus constraints and finite TTD range constraints in either far-field or near-field scenarios. Furthermore, the conventional TTD-based hybrid beamforming architecture employs a fully-connected structure. In such a structure, the number of TTDs and their required maximum delay are proportional to the aperture of the entire antenna array [15, 20]. Therefore, for ELAAs with large apertures, the requirement on TTDs can be stringent and costly. Motivated by this, we proposed a new sub-connected TTD-based architecture that not only alleviates the stringent TTD requirement but also enhance energy efficiency. Finally, considering that the far-field model is essentially an approximation of the near-field model [3], our study primarily focuses on beamforming design in near-field systems. Our design can be directly extended to far-field systems. The primary contributions of this paper can be succinctly summarized below, which are boldly and explicitly contrasted to the relevant state-of-the-art in Table I.
@@
  • We propose a penalty-based full digital approximation (FDA) method for near-field beamforming designs with full channel state information (CSI) in both fully-connected and sub-connected architectures. In this approach, the optimization variables are iteratively optimized by closed-form solutions or low-complexity one-dimensional search. This approach also guarantees convergence to the stationary point of the spectral efficiency maximization problem.
  • Then, we propose a low-complexity heuristic two-stage (HTS) near-field beamforming design method based on a low-complexity channel estimation protocol. Specifically, the analog beamformer is designed to maximize the line-of-sight (LoS) signal power for each user using partial CSI obtained from the low-complexity near-field beam training. To facilitate the analog beamforming design, a high-quality piecewise-near-field approximation method is proposed. Then, the digital beamformer is optimized to maximize the spectral efficiency through the low-dimensional equivalent channel.
  • Numerical results demonstrate the efficiency of the proposed method, revealing that 1) the spatial wideband effect can be effectively eliminated by utilizing TTD; 2) the penalty-based FDA method yields the best performance in both fully-connected and sub-connected architectures; 3) the HTS method can achieve comparable performance to the penalty-based FDA method in most cases; and 4) sub-connected architectures achieve higher energy efficiency and more robust performance than fully-connected architectures for different bandwidths and TTD maximum delays.
@@ I-C Organization and Notations
Notations: Scalars, vectors, and matrices are represented by the lower-case, bold-face lower-case, and bold-face upper-case letters, respectively; The transpose, conjugate transpose, pseudo-inverse, and trace of a matrix are denoted by $\mathbf{X}^{T}$, $\mathbf{X}^{H}$, $\mathbf{X}^{\dagger}$, and $\mathrm{tr}(\mathbf{X})$, respectively. The Euclidean norm of vector $\mathbf{x}$ is denoted as $\|\mathbf{x}\|$, while the Frobenius norm of matrix $\mathbf{X}$ is denoted as $\|\mathbf{X}\|_{F}$. For matrix $\mathbf{X}$, $\mathbf{X}(i,j)$ refers to its entry in the $i$-th row and $j$-th column. $\mathbf{X}(i:j,:)$ represents a matrix composed of the rows from the $i$-th to the $j$-th, and $\mathbf{X}(:,i)$ represents a vector composed of the $i$-th column. $\mathbf{x}(i)$ indicates the $i$-th entry of vector $\mathbf{x}$. A block diagonal matrix with diagonal blocks $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$ is denoted as $\mathrm{blkdiag}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})$. $\mathbb{E}[\cdot]$ represents the statistical expectation, while $\mathrm{Re}\{\cdot\}$ refers to the real component of a complex number. $\mathcal{CN}(\mu,\sigma^{2})$ denotes the circularly symmetric complex Gaussian random distribution with mean $\mu$ and variance $\sigma^{2}$. The matrix $\mathbf{I}_{N}$ denotes the identity matrix of size $N$. The symbol $\angle$ represents the phase of a complex value.
@@ II System Model and Problem Formulation
We consider a THz near-field wideband multi-user communication system in which a BS utilizes a uniform linear array (ULA) consisting of $N$ antenna elements with an antenna spacing of $d$. The system involves $K$ communication users, each equipped with a single antenna, and their indices are collected in the set $\mathcal{K}$. To mitigate inter-symbol interference inherent in wideband communications, orthogonal frequency division multiplexing (OFDM) is employed, where the signal is generated in the frequency domain and then transformed into the time domain by inverse fast Fourier transform (IFFT). Let $M$ and $L_{\mathrm{CP}}$ denote the number of subcarriers and the length of cyclic prefix in OFDM, respectively, $B$ denote the system bandwidth, $f_{c}$ denote the central carrier frequency, $c$ denote the speed of light, and $\lambda_{c}=c/f_{c}$ denote the wavelength of the central carrier. Consequently, the frequency of subcarrier $m$ is determined as $f_{m}=f_{c}+B(2m-1-M)/(2M),\forall m\in\mathcal{M}=\{1,...,M\}$. The antenna spacing of the ULA at the BS is set to half of the central wavelength, i.e., $d=\lambda_{c}/2$.
@@ II-A Near-Field Wideband Channel Model
We assume that all users are located in the near-field region of the BS, i.e., the distances between users and the BS are less than the Rayleigh distance $2D^{2}/\lambda_{c}$ [3], where $D=(N-1)d$ denotes the aperture of the antenna array. Therefore, as shown in Fig. 1, we adopt the multipath near-field channel model for each communication user, which consists of a LoS channel and several non-line-of-sight (NLoS) channels induced by $L_{k}$ scatterers associated with user $k$. Let $r_{k}$ and $\theta_{k}$ denote the distance and angle of user $k$ with respect to the central element of the ULA, respectively, and $\tilde{r}_{k,l}$ and $\tilde{\theta}_{k,l}$ denote the counterpart of the $l$-th scatterer for user $k$. Based on the uniform spherical wave model [3], the near-field channel between the BS and user $k$ at the $m$-th subcarrier can be modelled as
@@
Here, $\beta_{m,k}$ and $\tilde{\beta}_{m,k,l}$ denote the complex channel gain for LoS and NLoS channels, respectively. The channel gain for the LoS channel can be modelled as [23]
@@
We note that in THz bands, substantial reflection losses lead to NLoS channels experiencing attenuation greater than $10$ dB on average when compared to LoS channels [25]. Consequently, the communication channel is LoS-dominant and NLoS-assisted.
Furthermore, vector $\mathbf{b}(f,\theta,r)$ denotes the array response vector and is given by
@@
where $r_{n}(r,\theta)$ denote the propagation distance of the signal emitted from the $n$-th antenna. In the near-field region, it has to be accurately modelled as
@@ II-B Antenna Architectures for Wideband Beamfocusing
As highlighted in Section I, TTD-based hybrid beamforming architectures present a viable solution to address challenges arising from the frequency-dependent array response. Distinct from conventional PSs, TTDs offer the capability to achieve a frequency-dependent phase shift, denoted as $e^{-j2\pi f_{m}t}$, at subcarrier $m$ within the frequency domain by introducing a time delay $t$ into the signals [14]. However, the deployment of TTDs is often limited by their high cost and energy consumption. Consequently, TTD-based hybrid beamforming architectures typically incorporate only a limited number of TTDs between the RF chains and the PS networks. In the conventional fully-connected architecture, as illustrated in Fig. 2(a), each TTD is linked to a sub-array through PSs, and each RF chain is connected to the entire antenna array. To further improve energy efficiency and reduce TTD hardware requirements, we introduce a new sub-connected setup for TTD-based hybrid beamforming, depicted in Fig. 2(b), where each RF chain is connected to a sub-array via TTDs and PSs.
@@
Let $N_{\mathrm{RF}}$ denote the number of RF chains in the TTD-based hybrid beamforming architecture, which is subject to the constraint $K\leq N_{\mathrm{RF}}\ll N$, and $N_{\mathrm{T}}$ denote the number of TTDs connected to each RF chain. We assume that each TTD is connected to a sub-array with the same size. Then, when the TTD-hybrid beamforming architecture is exploited, the received signal of user $k$ at subcarrier $m$ can be modelled as
@@
where matrices $\mathbf{A}\in\mathbb{C}^{N\times N_{\mathrm{T}}N_{\mathrm{RF}}}$ and $\mathbf{T}_{m}\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{RF}}\times N_{\mathrm{RF}}}$ denote the frequency-independent and frequency-dependent analog beamformers realized by TTDs and PSs, respectively, matrix $\mathbf{D}_{m}\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$ denotes the baseband digital beamformer, vector $\mathbf{s}_{m}\in\mathbb{C}^{K\times 1}$ denotes the information symbols of the $K$ users at subcarrier $m$, satisfying $\mathbb{E}[\mathbf{s}_{m}\mathbf{s}_{m}^{H}]=\mathbf{I}_{K}$, and $n_{m,k}\sim\mathcal{CN}(0,\sigma_{m,k}^{2})$ represents the additive Gaussian white noise. In particular, for both fully-connected and sub-connected architectures, the TTD-based analog beamformer is given by
@@
where $\mathbf{t}_{n}=[t_{n,1},\dots,t_{n,N_{\mathrm{T}}}]\in\mathbb{R}^{N_{\mathrm{T}}\times 1}$ denotes the time-delay vector realized the TTDs connected to the $n$-th RF chain. The time delay of each TTD needs to satisfy the maximum delay constraint, i.e., $t_{n,l}\in[0,t_{\max}],\forall n=1,...,N_{\mathrm{RF}},l=1,\dots,N_{\mathrm{T}}$. The PS-based analog beamformer $\mathbf{A}$ has different structures for the fully-connected and sub-connected architectures, which are detailed in the following.
@@
    where $\mathbf{A}_{n}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{T}}}$ and $\mathbf{a}_{n,l}\in\mathbb{C}^{\frac{N_{\mathrm{sub}}}{N_{\mathrm{T}}}\times 1}$ denote the PS-based analog beamformer for the sub-array connected to the $n$-th RF chain and the $l$-th TTD of this chain, respectively.
@@
Compared to the fully-connected architecture, the advantages of the sub-connected architecture are summarized as follows. Firstly, it significantly reduces the required number of PSs from $NN_{\mathrm{RF}}$ PSs in the fully-connected architecture to just $N$ PSs in the sub-connected architecture. Secondly, in the sub-connected architecture, the TTDs connected to each RF chain only need to mitigate the spatial wideband effect of individual sub-arrays. It is worth noting that this effect is relatively less significant compared to the effect related to the entire array. Consequently, only a few TTDs are required for each RF chain and the maximum delay requirement of TTDs is also significantly reduced.

### v6

@@ Abstract
A near-field wideband communication system is investigated in which a base station (BS) employs an extra-large scale antenna array (ELAA) to serve multiple users in its near-field region. To facilitate near-field multi-user beamforming and mitigate the spatial wideband effect, the BS employs a hybrid beamforming architecture based on true-time delayers (TTDs). In addition to the conventional fully-connected TTD-based hybrid beamforming architecture, a new sub-connected architecture is proposed to improve energy efficiency and reduce hardware requirements. Two wideband beamforming optimization approaches are proposed to maximize spectral efficiency for both architectures. 1) Fully-digital approximation (FDA) approach: In this method, the TTD-based hybrid beamformer is optimized by the block-coordinate descent and penalty method to approximate the optimal digital beamformer. This approach ensures convergence to the stationary point of the spectral efficiency maximization problem. 2) Heuristic two-stage (HTS) approach: In this approach, the analog and digital beamformers are designed in two stages. In particular, two low-complexity methods are proposed to design the high-dimensional analog beamformers based on approximate and exact line-of-sight channels, respectively. Subsequently, the low-dimensional digital beamformer is optimized based on the low-dimensional equivalent channels, resulting in reduced computational complexity and channel estimation complexity. Our numerical results show that 1) the proposed approach effectively eliminates the spatial wideband effect, and 2) the proposed sub-connected architecture is more energy efficient and has fewer hardware constraints on the TTD and system bandwidth compared to the fully-connected architecture.
@@
Beamfocusing, near-field, spatial wideband effect, true-time delayers, wideband communications
@@ I Introduction
     Near-field effect     ✔     ✘     ✘     ✔     ✔     ✔
     Spatial wideband effect     ✘     ✔     ✔     ✔     ✔     ✔
     Multiple communication users     ✔     ✘     ✘     ✘     ✘     ✔
     The employment of TTDs     ✘     ✘     ✔     ✘     ✔     ✔
@@ I-B Motivations and Contributions
Against the above background, the TTD-based hybrid beamforming architecture is a promising solution for the wideband beamforming design in either far-field or near-field systems. However, it is notable that most existing research in this area has predominantly focused on single-user systems [14, 16, 17, 23, 18, 19, 21, 22]. In such contexts, beamforming design is typically oriented towards maximizing array gain along the direction or location of the user. This approach, while effective for single-user scenarios, is flawed in multi-user systems where both array gain and inter-user interference must be carefully balanced. This requirement introduces a complex design challenge, particularly in the context of TTD-based hybrid beamforming. To the best of the authors’ knowledge, this is the first attempt to directly address the design problem of TTD-based hybrid beamforming in multiuser systems by considering both unit modulus constraints and finite TTD range constraints in either far-field or near-field scenarios. Furthermore, the conventional TTD-based hybrid beamforming architecture employs a fully-connected structure. In such a structure, the number of TTDs and their required maximum delay are proportional to the aperture of the entire antenna array [16, 21]. Therefore, for ELAAs with large apertures, the requirement on TTDs can be stringent and costly. Motivated by this, we proposed a new sub-connected TTD-based architecture that not only alleviates the stringent TTD requirement but also enhances energy efficiency. Finally, considering that the far-field model is essentially an approximation of the near-field model [4], our study primarily focuses on beamforming design in near-field systems. Our design can be directly extended to far-field systems. The primary contributions of this paper can be succinctly summarized below, which are boldly and explicitly contrasted to the relevant state-of-the-art in Table I.
@@
  • We propose a penalty-based fully-digital approximation (FDA) method for near-field beamforming designs in both fully-connected and sub-connected architectures. In this approach, the optimization variables are iteratively optimized by closed-form solutions or low-complexity one-dimensional search. This approach also guarantees convergence to the stationary point of the spectral efficiency maximization problem.
  • We then propose a low-complexity heuristic two-stage (HTS) near-field beamforming design method. In particular, the analog beamformer is first designed to maximize the line-of-sight (LoS) signal power for each user. We propose two low-complexity methods to design the analog beamformers based on the approximated and exact LoS channels, respectively. Subsequently, the digital beamformer is optimized to maximize spectral efficiency through the low-dimensional equivalent channel.
  • Numerical results demonstrate the efficiency of the proposed method, revealing that 1) the spatial wideband effect can be effectively eliminated by utilizing TTD; 2) the penalty-based FDA method yields the best performance in both fully-connected and sub-connected architectures; 3) the HTS method can achieve comparable performance to the penalty-based FDA method in most cases; and 4) Compared to fully-connected architectures, sub-connected architectures demonstrate superior energy efficiency and enhanced robustness in the face of ultra-large bandwidths and low TTD maximum delays.
@@ I-C Organization and Notations
Notations: Scalars, vectors, and matrices are represented by the lower-case, bold-face lower-case, and bold-face upper-case letters, respectively; The transpose, conjugate transpose, pseudo-inverse, and trace of a matrix are denoted by $\mathbf{X}^{T}$, $\mathbf{X}^{H}$, $\mathbf{X}^{\dagger}$, and $\mathrm{tr}(\mathbf{X})$, respectively. The Euclidean norm of vector $\mathbf{x}$ is denoted as $\|\mathbf{x}\|$, while the Frobenius norm of matrix $\mathbf{X}$ is denoted as $\|\mathbf{X}\|_{F}$. For matrix $\mathbf{X}$, $\mathbf{X}(i,j)$ refers to its entry in the $i$-th row and $j$-th column. $\mathbf{X}(i:j,:)$ represents a matrix composed of the rows from the $i$-th to the $j$-th, and $\mathbf{X}(:,i)$ represents a vector composed of the $i$-th column. $\mathbf{x}(i)$ indicates the $i$-th entry of vector $\mathbf{x}$. A block diagonal matrix with diagonal blocks $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$ is denoted as $\mathrm{blkdiag}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})$. $\mathbb{E}[\cdot]$ represents the statistical expectation, while ${}\Re\{\cdot\}$ refers to the real component of a complex number. $\mathcal{CN}(\mu,\sigma^{2})$ denotes the circularly symmetric complex Gaussian random distribution with mean $\mu$ and variance $\sigma^{2}$. The matrix $\mathbf{I}_{N}$ denotes the identity matrix of size $N$. The symbol $\angle$ represents the phase of a complex value.
@@ II System Model and Problem Formulation
We consider a THz near-field wideband multi-user communication system in which a BS utilizes a uniform linear array (ULA) consisting of $N$ antenna elements with an antenna spacing of $d$. The system involves $K$ communication users, each equipped with a single antenna, and their indices are collected in the set $\mathcal{K}$. To mitigate inter-symbol interference inherent in wideband communications, orthogonal frequency division multiplexing (OFDM) is employed, where the signal is generated in the frequency domain and then transformed into the time domain by inverse discrete Fourier transform (IDFT). Let $M$ and $L_{\mathrm{CP}}$ denote the number of subcarriers and the length of cyclic prefix in OFDM, respectively, $B$ denote the system bandwidth, $f_{c}$ denote the central carrier frequency, $c$ denote the speed of light, and $\lambda_{c}=c/f_{c}$ denote the wavelength of the central carrier. Consequently, the frequency of subcarrier $m$ is determined as $f_{m}=f_{c}+B(2m-1-M)/(2M),\forall m\in\mathcal{M}=\{1,...,M\}$. The antenna spacing of the ULA at the BS is set to half of the central wavelength, i.e., $d=\lambda_{c}/2$.
@@ II-A Near-Field Wideband Channel Model
We assume that all users are located in the near-field region of the BS, i.e., the distances between users and the BS are less than the Rayleigh distance $2D^{2}/\lambda_{c}$ [4], where $D=(N-1)d$ denotes the aperture of the antenna array. Therefore, as shown in Fig. 1, we adopt the multipath near-field channel model for each communication user, which consists of a LoS channel and several non-line-of-sight (NLoS) channels induced by $L_{k}$ scatterers associated with user $k$. Let $r_{k}$ and $\theta_{k}$ denote the distance and angle of user $k$ with respect to the central element of the ULA, respectively, and $\tilde{r}_{k,l}$ and $\tilde{\theta}_{k,l}$ denote the counterpart of the $l$-th scatterer for user $k$. Based on the uniform spherical wave model [4], the channel between the BS and user $k$ at the $m$-th subcarrier can be modeled as [^1] Here, we assume that the discrete Fourier transform (DFT) operation and cyclic prefix removal have been conducted at the OFDM receiver. We refer to [9] and [24] for the detailed derivation of the frequency-domain channels in OFDM systems.
@@
Here, $\beta_{m,k}$ and $\tilde{\beta}_{m,k,l}$ denote the complex channel gain for LoS and NLoS channels, respectively. The channel gain for the LoS channel can be modeled as [25]
@@
We note that in THz bands, substantial reflection losses lead to NLoS channels experiencing attenuation greater than $10$ dB on average when compared to LoS channels [27]. Consequently, the communication channel is LoS-dominant and NLoS-assisted. Furthermore, vector $\mathbf{b}(f,\theta,r)$ denotes the array response vector and is given by
@@
where $r_{n}(r,\theta)$ denote the propagation distance of the signal emitted from the $n$-th antenna. In the near-field region, it has to be accurately modeled as
@@ II-B Antenna Architectures for Wideband Beamfocusing
As highlighted in Section I, TTD-based hybrid beamforming architectures present a viable solution to address challenges arising from the frequency-dependent array response. Unlike conventional PSs [^2] Due to the physical properties of the components used in implementing PSs (like capacitors and inductors), which have frequency-dependent behaviors, the response of PSs naturally changes with frequencies. However, unlike TTDs, the frequency-dependent behavior of PSs is fixed according to the circuit design and cannot be dynamically controlled. Such a property may complicate the system design. Therefore, in wideband systems, the PSs are usually designed to provide an approximately flat response across the entire working frequency band [28, 29]. Therefore, in this paper, we assume that the response of PSs is perfectly frequency-flat as in the literature [14, 16]., TTDs can realize a frequency-dependent phase shift of $e^{-j2\pi f_{m}t}$ on subcarrier $m$ in the frequency domain by introducing a time delay $t$ into the signals [15]. However, due to the high hardware cost and energy consumption of TTDs, TTD-based hybrid beamforming architectures usually install only a limited number of TTDs between the RF chain and the PS network [14, 16]. In the conventional fully-connected architecture, as illustrated in Fig. 2(a), each TTD is linked to a sub-array through PSs and each RF chain is connected to the entire antenna array. To further improve energy efficiency and reduce TTD hardware requirements, we introduce a new sub-connected setup for TTD-based hybrid beamforming, depicted in Fig. 2(b), where each RF chain is connected to a sub-array via TTDs and PSs.
@@
Let $N_{\mathrm{RF}}$ denote the number of RF chains in the TTD-based hybrid beamforming architecture, which is subject to the constraint $K\leq N_{\mathrm{RF}}\ll N$, and $N_{\mathrm{T}}$ denote the number of TTDs connected to each RF chain. We assume that each TTD is connected to a sub-array with the same size. Then, when the TTD-hybrid beamforming architecture is exploited, the received signal of user $k$ on subcarrier $m$ is given by
@@
where matrices $\mathbf{A}\in\mathbb{C}^{N\times N_{\mathrm{T}}N_{\mathrm{RF}}}$ and $\mathbf{T}_{m}\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{RF}}\times N_{\mathrm{RF}}}$ denote the frequency-independent and frequency-dependent analog beamformers realized by TTDs and PSs, respectively, matrix $\mathbf{D}_{m}\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$ denotes the baseband digital beamformer, vector $\mathbf{s}_{m}\in\mathbb{C}^{K\times 1}$ denotes the information symbols of the $K$ users on subcarrier $m$, satisfying $\mathbb{E}[\mathbf{s}_{m}\mathbf{s}_{m}^{H}]=\mathbf{I}_{K}$, and $n_{m,k}\sim\mathcal{CN}(0,\sigma_{m,k}^{2})$ represents the additive Gaussian white noise. For both fully-connected and sub-connected architectures, the TTD-based analog beamformer is given by
@@
where $\mathbf{t}_{n}=[t_{n,1},\dots,t_{n,N_{\mathrm{T}}}]^{T}\in\mathbb{R}^{N_{\mathrm{T}}\times 1}$ denotes the time-delay vector realized the TTDs connected to the $n$-th RF chain. The time delay of each TTD needs to satisfy the maximum delay constraint, i.e., $t_{n,l}\in[0,t_{\max}],\forall n=1,...,N_{\mathrm{RF}},l=1,\dots,N_{\mathrm{T}}$. Typically, achieving a larger $t_{\max}$ requires a more complex structure and a larger footprint size of TTDs [19, 30]. The PS-based analog beamformer $\mathbf{A}$ has different structures for the fully-connected and sub-connected architectures, which are detailed in the following.
@@
    where $\mathbf{A}_{n}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{T}}}$ and $\mathbf{a}_{n,l}\in\mathbb{C}^{\frac{N_{\mathrm{sub}}}{N_{\mathrm{T}}}\times 1}$ denote the PS-based analog beamformer for the sub-array connected to the $n$-th RF chain and its $l$-th TTD, respectively.
@@
Compared to fully-connected architectures, sub-connected architectures have the following advantages. First, it significantly reduces the required number of PSs from $NN_{\mathrm{RF}}$ to $N$. Second, in the sub-connected architecture, the TTDs connected to each RF chain only need to mitigate the spatial wideband effect of individual sub-arrays, which is less significant compared to the effect related to the entire array. Therefore, the requirements on the number of TTDs and the maximum delay are more relaxed.